A liquid of density 1390 kg/m3 flows steadily through a pipe of varying diameter and height. At Location 1 along the pipe, the flow speed is 9.45 m/s and the pipe diameter d1 is 10.3 cm. At Location 2, the pipe diameter d2 is 14.9 cm. At Location 1, the pipe is Δy=9.91 m higher than it is at Location 2. Ignoring viscosity, calculate the difference ΔP between the fluid pressure at Location 2 and the fluid pressure at Location 1.
Answer #1

Gravitational acceleration = g = 9.81 m/s2

Density of liquid = \rho = 1390 kg/m3

Pressure of the liquid at point 1 = P1

Diameter of pipe at point 1 = D1 = 10.3 cm = 0.103 m

Cross-sectional area of pipe at point 1 = A1 = \piD12/4

Speed of flow of liquid at point 1 = V1 = 9.45 m/s

Height of point 1 compared to point 2 = Y = 9.91 m

Pressure of the liquid at point 2 = P2

Diameter of pipe at point 2 = D2 = 14.9 cm = 0.149 m

Cross-sectional area of pipe at point 2 = A2 = \piD22/4

Speed of flow of liquid at point 2 = V2

Difference of fluid pressure between point 2 and point 1 = \DeltaP = P2 - P1

By continuity equation,

A1V1 = A2V2

(\piD12/4)V1 = (\piD22/4)V2

D12V1 = D22V2

(0.103)2(9.45) = (0.149)2V2

V2 = 4.516 m/s

By bernoulli's equation,

+ਬ = {60 + +u

P2 - R = ppp+ PgY - para

AP = P(V! - vit +PgY

ΔΡ = o (1390) ((9.45)2 – (45162) + (1390) (9.81)(9.91)

\DeltaP = 183023 Pa

Difference between the fluid pressure at location 2 and location 1 = 183023 Pa